OpenStudy (anonymous):

write an equation in slope-intercept form of the line with a slope of -3 and contains the point (4, -5)

OpenStudy (aroub):

y=mx+b , m=-3 , (4,-5) -5=-3(4)+b -5=-12+b 7=b y=-3x+7